OpenStudy (saiyan):

9. what is the simplest form of the equation? square root of 2 + square root of 5/square root of 2 - square root of 5.

OpenStudy (saiyan):

Anyone help plz.

OpenStudy (sohailiftikhar):

Do mean that \[\sqrt{2}+\sqrt{5}\div \sqrt{2}-\sqrt{5}\]

OpenStudy (saiyan):

yes

OpenStudy (sohailiftikhar):

ok

OpenStudy (saiyan):

I just don't know the steps to work through it.

OpenStudy (sohailiftikhar):

Do you want factorize it ?

OpenStudy (saiyan):

do u mean factor it

OpenStudy (sohailiftikhar):

you can multiply it with \[\sqrt{2}.\sqrt{5}\]

OpenStudy (sohailiftikhar):

to make it simple

OpenStudy (saiyan):

\[so \sqrt{2} \times \sqrt{5} which is \sqrt{10}\]

OpenStudy (saiyan):

or do i multiply the equation times sqrt(5)

OpenStudy (danjs):

Is it supposed to be.. \[\frac{ \sqrt{2} + \sqrt{5} }{ \sqrt{2} -\sqrt{5} }\]

OpenStudy (saiyan):

yes

OpenStudy (sohailiftikhar):

\[dude when you will multiply then \sqrt{2}*\sqrt{2} will become 2 and vice versa\]

OpenStudy (danjs):

multiply the fraction by.. \[\frac{ \sqrt{2} + \sqrt{5} }{ \sqrt{2} + \sqrt{5} }\]

OpenStudy (danjs):

\[(\frac{ \sqrt{2} + \sqrt{5} }{ \sqrt{2} -\sqrt{5} }) * (\frac{ \sqrt{2} + \sqrt{5} }{ \sqrt{2} +\sqrt{5} }) =\]

OpenStudy (danjs):

no more roots in the denominator after that

OpenStudy (sohailiftikhar):

yes

OpenStudy (danjs):

\[=\frac{ 2 + 2\sqrt{2}\sqrt{5} + 5 }{ 2 - 5 }\]

OpenStudy (danjs):

\[=\frac{ 7 + 2\sqrt{2}\sqrt{5} }{ -3 }\] that about it

OpenStudy (sohailiftikhar):

yeah!

OpenStudy (saiyan):

so you just multiply the whole equation by the numerator

OpenStudy (danjs):

switch the sign in the denominator

OpenStudy (danjs):

+ to -, or, - to +

OpenStudy (sohailiftikhar):

yes you can but you have to multiply the both sides

OpenStudy (danjs):

you are just multiplying the fraction by 1

OpenStudy (danjs):

\[\frac{ \sqrt{2} + \sqrt{5} }{ \sqrt{2} + \sqrt{5} }=1\]

OpenStudy (sohailiftikhar):

if you want to remove the numerator then multiply the both sides of equation with it

OpenStudy (saiyan):

\[=7+2\sqrt{2}√5/−3\] is the answer

OpenStudy (sohailiftikhar):

yes

OpenStudy (saiyan):

ok thx guys

OpenStudy (sohailiftikhar):

no problem dude